OpenStudy (anonymous):

How would i put this in standard form? Standard Form- y=ax^2 + bx +c y = (x - 7)² + 24

OpenStudy (anonymous):

and how would i graph these? y = -0.1(x - 7)² + 24 y = 4(x - 7)² + 24

OpenStudy (anonymous):

@hartnn